Please find the equation of a line perpendicular to y=2x/5=3 with the y-intercept being (0,1)
Answer by Cromlix(4381)

Lines that are perpendicular to one another
have slopes (gradients) that multiply together
to give -1
m1 * m2 = -1
Assuming your equation is:
y = 2/5x - 3
Gradient (slope) = 2/5
m1 * m2 = -1
2/5 * m2 = -1
m2 = -5/2
Using equation:
y = mx + c
m = -5/2 and c = 1
y = -5/2x + 1
Hope this helps.
:-)